OpenStudy (anonymous):

Choose the equation that best represents the general form of the following expression. x^2/2=7x - x 2 + 14x + 0 x 2 + 14x = 0 x 2 - 14x = 0 x 2 - 7x + 2 = 0

OpenStudy (anonymous):

x^2 = 14x x^2 - 14x = 0.

OpenStudy (anonymous):

the general form is ax^2 +bx +c = 0 In this case c=0